The consumption of chicken changes from 10kg per month to 5kg per month due to salary increase of R500 from R2000 initially received monthly. What is the income elasticity of the demand for chicken?


1
Expert's answer
2022-04-15T16:38:48-0400

"IED=\frac{10-5}{10+5}\cdot\frac{500+2000}{500-2000}=-0 .56."